Docosanol is the generic name for the drug Abreva which is used to treat cold sores. It
was found to be 80.90% carbon, 4.899% oxygen, and the rest was hydrogen. If the
molecular mass of docosanol is 326.6 amu, what is the molecular formula of docosanol?
1
Expert's answer
2019-11-06T07:45:56-0500

W(C) = 0,809

W(O) = 0,04899

W(H) = 0,14201

M = 326.6 amu

n(C) = M*W(C)/M(C) = 326,6*0,809/12 = 22

n(O) = M*W(O)/M(O) = 326,6*0,04899/16 = 1

n(H) = M*W(H)/M(H) = 326,6*0,14201/1 = 46

The molecular formula of docosanol is C22H46O